THE INFRARED SPECTRUM OF CH_{2}^{+} REVISITED

\underline{CHRISTOPHER F. NEESE} and TAKESHI OKA, Department of Chemistry, Department of Astronomy & Astrophysics, and the Enrico Fermi Institute, The University of Chicago, Chicago IL, 60637, USA.

The first reported spectrum of CH_{2}^{+} was the K_{a} = 0 series of the ν_{3} fundamental, recorded by our group in 1991.^{a} Subsequently, the K_{a} = 1 series was observed in 1993 and later analyzed by Jensen et al.^{b} These spectra were recorded in liquid-nitrogen-cooled positive-column discharges of CH_{4}, H_{2}, and He. In the 1993 spectrum, absorption lines from CH_{2}^{+} were discriminated from those of CH_{3}^{+} by recording the spectrum of a discharge without H_{2}. In the absence of H_{2}, the CH_{2}^{+} lines grew stronger and the CH_{3}^{+} lines weakened. In the 1993 spectrum, weak lines attributed to CH_{2}^{+} with K_{a} = 2 were observed. The K_{a} = 2 lines were significantly weaker than the K_{a} = 1 lines because of a reduced Boltzman factor (A ≈ 67 cm^{−1}). The assignment of these lines was stalled because of their low intensity, the complicated spectral pattern of CH_{2}^{+} (due to the combined effects of quasi-linearity and the Renner-Teller effect), and interfering absorption lines of CH_{3}^{+}, C_{2}H_{2}^{+}, and C_{2}H_{3}^{+}.

Two new developments have led us to revisit the infrared spectrum of CH_{2}^{+}, with the goal of assigning K_{a} ≥ 2 lines. The first development is the observation of the Ã(0, 3, 0)^{3} ← X̃(0, 0, 0)^{2} electronic band of CH_{2}^{+}.^{c}^{,}^{d} This band provides combination differences that can be directly compared with the ν_{3} fundamental. The second development is the discovery of a new plasma chemistry — allene and He in a water-cooled positive-column discharge — that appears to be selective for CH_{2}^{+}. This new chemistry allows us to confirm CH_{2}^{+} as the carrier of previously observed lines. This paper will discuss progress made in the assignment of the infrared spectrum of CH_{2}^{+} in light of these new developments.

aM. Rösslein, C. M. Gabrys, M.-F. Jagod, and T. Oka, J. Mol. Spectrosc. 153, 738 (1992).

bP. Jensen, M. Brumm, W. P. Kraemer, and P. R. BunkerJ. Mol. Spectrosc. 172, 194 (1995).

cJ. L. Gottfried and T. Oka, J. Chem. Phys. 121, 11527 (2004).

dSee talk entitled “The Near-Infrared Spectrum of CH_{2}^{+}.”